# Complexity Results in Team Semantics: Nonemptiness Is Not So Complex

### v1

@@ 1 Introduction
The logic $\mathrm{PL}(\textsc{ne})$ is also interesting due to its team-semantic closure properties. Unlike dependence logic, $\mathrm{PL}(\textsc{ne})$ is not downward closed (where a logic is downward closed if for each of its formulas $\varphi$, the truth of $\varphi$ in a team $t$ implies truth in all subteams of $t$). Like inclusion logicm, it is union closed (the truth of a formula in a nonempty collection of teams implies its truth in the union of the collection); however, $\mathrm{PL}(\textsc{ne})$ differs from inclusion logic in that it does not have the empty team property (a logic has this property if all formulas are true in the empty team) and in that it is convex (the truth of a formula in two teams implies its truth in all teams between the two with respect to the subset relation). The convexity property has not been studied in the literature on team semantics until recently [24, 9, 6], and, to our knowledge, the literature does not contain any results on the complexity-theoretic properties of logics which are convex but not downward closed (convexity can be seen as a generalization of downward closure to a setting in which the empty team property fails).
@@
Table 1 lists our results alongside those for classical propositional logic $\mathrm{PL}$, propositional dependence logic $\mathrm{PL}(=\hskip-2.84544pt(\cdot))$ [39], propositional inclusion logic $\mathrm{PL}(\subseteq)$}, and propositional team logic $\mathrm{PL}(\sim)$ [36, 34, 33]. The logic $\mathrm{PL}(\sim)$ is expressively complete for all propositional team properties [30] and many of the common team-semantic connectives are definable in $\mathrm{PL}(\sim)$, so it provides a natural upper bound for the complexity of team-based extensions of $\mathrm{PL}$. The proof in [32] that $\mathrm{SAT}[\mathrm{PL}(=\hskip-2.84544pt(\cdot))]$ is NP-complete—i.e., that adding dependence atoms to $\mathrm{PL}$ does not increase the complexity of the satisfiability problem—makes use of downward closure. Similarly, the proof in [19] that $\mathrm{VAL}[\mathrm{PL}(\subseteq)]$ is coNP-complete—i.e., that adding inclusion atoms to $\mathrm{PL}$ does not increase the complexity of the validity problem—makes use of union closure; as does the proof of the P-completeness of $\mathrm{MC}[\mathrm{PL}(\subseteq)]$ in [21]. The complexity results for the logic $\mathrm{PL}(\textsc{ne})$ follow this pattern: being convex (downward closed up to a point), the complexity of its satisfiability problem is the same as that of $\mathrm{PL}$; being union closed, the complexity of its validity problem is the same as that of $\mathrm{PL}$, and its model-checking problem is tractable (our proofs of these results also make use of the relevant closure properties). As a result, we see that, at least with regard to the satisfiability and validity problems, adding the nonemptiness atom ne to $\mathrm{PL}$ does not increase complexity.
@@ 2 Preliminaries
where $p\in\mathsf{Prop}$. The language of $\mathrm{PL}(\textsc{ne})$ extends that of $\mathrm{PL}$ with the nonemptiness atom ne. [^1] Some formulations of the syntax of $\mathrm{PL}(\textsc{ne})$ in the literature feature a negation operator that can be applied to all $\mathrm{PL}$-formulas, with the semantics $t\models\alpha$ iff for all $v\in t$, $\{v\}\not\models\alpha$. Our results also hold for this variant of $\mathrm{PL}(\textsc{ne})$—this follows from the fact that formulas in this variant can be converted into negation normal form in polynomial time.
We let $\bigvee\emptyset:=\bot$, and $\bigwedge\emptyset:=\top$. We also call formulas of $\mathrm{PL}$ classical formulas. The metavariable $\alpha$ ranges over classical formulas, and $\Lambda$, over sets of classical formulas; metavariables denoted by other Greek letters range over all (sets of) $\mathrm{PL}(\textsc{ne})$-formulas. We write $\mathsf{P}(\varphi)$ for the set of propositional variables appearing in $\varphi$. We write $\varphi(\psi/\chi)$ for the result of replacing all occurrences of $\chi$ in $\varphi$ by $\psi$. We call formulas of the form $p$, $\lnot p$, $\bot$, $\top$, or ne literals.
@@
For any formula $\varphi$ and any team $t$ with domain $\supseteq\mathsf{P}(\varphi)$, the satisfaction relation $t\models\varphi$ is defined inductively by:
@@
For any set $\Gamma\cup\{\varphi\}$ of formulas, we write $t\models\Gamma$ if $t\models\psi$ for all $\psi\in\Gamma$, and we write $\Gamma\models\varphi$ if $t\models\Gamma$ implies $t\models\varphi$. We write $\varphi\models\psi$ for $\{\varphi\}\models\psi$. We write $\models\varphi$ if $t\models\varphi$ for all teams $t$ with domain $\supseteq\mathsf{P}(\varphi)$.
@@
convexity property iff $[t\models\varphi,u\models\varphi\text{ and }u\subseteq s\subseteq t]\implies s\models\varphi.$
It is easy to see that $\varphi$ is flat iff it has the empty team, downwards-closure, and union-closure properties.
@@
The right-to-left direction follows by Proposition 1. For the left-to-right direction, an easy induction on the complexity of $\varphi$ shows that if ne occurs in $\varphi$, then $\varphi$ does not have the empty team property. We only give the disjunction case here. If $\varphi=\psi\vee\chi$ and ne occurs in $\varphi$, then it must occur in $\psi$ or $\chi$. If it occurs in $\psi$, then, by the induction hypothesis, $\psi$ does not have the empty team property, but then clearly neither does $\psi\vee\chi$. Similarly if ne occurs in $\chi$.
@@ 3 Model-checking Problem
Let $\mathrm{Maxsub}(t,\varphi)=\mathrm{maxsub}(t,\varphi)$ if $\mathrm{maxsub}(t,\varphi)$ exists, and $\mathrm{Maxsub}(t,\varphi)=\mathbf{0}$ otherwise, where $\mathbf{0}$ is a “null team” such that $\mathbf{0}\not\models\varphi$ for all $\varphi$. Let $|t|$ denote the size of $t$, and $|\varphi|$ the number of symbols in $\varphi$.
@@
First check whether $\varphi=\textsc{ne}$. If yes, then if $t=\emptyset$, return $\mathbf{0}$, and otherwise return $t$. Clearly this can be done in polynomial time with respect to $|t|+|\varphi|$. If $\varphi\neq\textsc{ne}$, check $\mathrm{maxsub}(t,\varphi)$; the result then follows by [21, Lemma 5.1].
@@
A function $f:\mathrm{subOcc}(\varphi)\to 2^{\mathsf{P(\varphi)}}\cup\{\mathbf{0}\}$ is called a labelling function. We describe an algorithm for computing a sequence $f_{0},f_{1},\ldots$ of such labelling functions. In what follows, we let $\mathbf{0}\cap s:=\mathbf{0}$ and $\mathbf{0}\cup s:=\mathbf{0}$ for all teams $s$.
@@ 5 Validity Problem
We make use of the following notion: the flattening $\varphi^{f}\in\mathrm{PL}$ of $\varphi\in\mathrm{PL}(\textsc{ne})$ is defined by $\varphi^{f}:=\varphi(\top/\textsc{ne})$ (see [5]; c.f. [36]). It is easy to see that $\varphi\models\varphi^{f}$; observe also that by Proposition 2, $\varphi\in\mathrm{PL}(\textsc{ne})$ has the empty team property iff $\varphi=\varphi^{f}$.
@@ 6.0.1 Acknowledgements
The first author received support from the Academy of Finland, decision number 368671, and from the European Research Council (ERC), grant 101020762. The authors would like to thank Maria Aloni for helpful discussions related to the content of this article.

### v2

@@ 1 Introduction
The logic $\mathrm{PL}(\textsc{ne})$ is also interesting due to its team-semantic closure properties. Unlike dependence logic, $\mathrm{PL}(\textsc{ne})$ is not downward closed (where a logic is downward closed if for each of its formulas $\varphi$, the truth of $\varphi$ in a team $t$ implies truth in all subteams of $t$). Like inclusion logic, it is union closed (the truth of a formula in a nonempty collection of teams implies its truth in the union of the collection); however, $\mathrm{PL}(\textsc{ne})$ differs from inclusion logic in that it does not have the empty team property (a logic has this property if all formulas are true in the empty team) and in that it is convex (the truth of a formula in two teams implies its truth in all teams between the two with respect to the subset relation). The convexity property has not been studied in the literature on team semantics until recently [25, 10, 5], and, to our knowledge, the literature does not contain any results on the complexity-theoretic properties of logics which are convex but not downward closed (convexity can be seen as a generalization of downward closure to a setting in which the empty team property fails).
@@
Table 1 lists our results alongside those for classical propositional logic $\mathrm{PL}$, propositional dependence logic $\mathrm{PL}(=\hskip-2.84544pt(\cdot))$ [41], propositional inclusion logic $\mathrm{PL}(\subseteq)$}, and propositional team logic $\mathrm{PL}(\sim)$ [36, 35, 34]. The logic $\mathrm{PL}(\sim)$ is expressively complete for all propositional team properties [31] and many of the common team-semantic connectives are definable in $\mathrm{PL}(\sim)$, so it provides a natural upper bound for the complexity of team-based extensions of $\mathrm{PL}$. The proof in [33] that $\mathrm{SAT}[\mathrm{PL}(=\hskip-2.84544pt(\cdot))]$ is NP-complete—i.e., that adding dependence atoms to $\mathrm{PL}$ does not increase the complexity of the satisfiability problem—makes use of downward closure. Similarly, the proof in [19] that $\mathrm{VAL}[\mathrm{PL}(\subseteq)]$ is coNP-complete—i.e., that adding inclusion atoms to $\mathrm{PL}$ does not increase the complexity of the validity problem—makes use of union closure; as does the proof of the P-completeness of $\mathrm{MC}[\mathrm{PL}(\subseteq)]$ in [22]. The complexity results for the logic $\mathrm{PL}(\textsc{ne})$ follow this pattern: being convex (downward closed up to a point), the complexity of its satisfiability problem is the same as that of $\mathrm{PL}$; being union closed, the complexity of its validity problem is the same as that of $\mathrm{PL}$, and its model-checking problem is tractable (our proofs of these results also make use of the relevant closure properties). As a result, we see that, at least with regard to the satisfiability and validity problems, adding the nonemptiness atom ne to $\mathrm{PL}$ does not increase complexity—in contrast, and as might be inferred from Table 1, many other team-based atoms and connectives do increase complexity, often quite considerably.
@@ 2 Preliminaries
where $p\in\mathsf{Prop}$. The language of $\mathrm{PL}(\textsc{ne})$ extends that of $\mathrm{PL}$ with the nonemptiness atom ne. [^1] Some formulations of the syntax of $\mathrm{PL}(\textsc{ne})$ in the literature feature a negation operator $\lnot$ that can be applied to all $\mathrm{PL}$-formulas, with the semantics $t\models\lnot\alpha$ iff for all $v\in t$, $\{v\}\not\models\alpha$. Our results also hold for this variant of $\mathrm{PL}(\textsc{ne})$—this follows from the fact that formulas in this variant can be converted into negation normal form in polynomial time.
We let $\bigvee\emptyset:=\bot$, and $\bigwedge\emptyset:=\top$. We also call formulas of $\mathrm{PL}$ classical formulas. The metavariable $\alpha$ ranges over classical formulas; and $\Lambda$, over sets of classical formulas; metavariables denoted by other Greek letters range over all (sets of) $\mathrm{PL}(\textsc{ne})$-formulas. We write $\mathsf{P}(\varphi)$ for the set of propositional variables appearing in $\varphi$. We write $\varphi(\psi/\chi)$ for the result of replacing all occurrences of $\chi$ in $\varphi$ by $\psi$. We call formulas of the form $p$, $\lnot p$, $\bot$, $\top$, or ne literals.
@@
For any formula $\varphi$ and any team $t$ with domain containing $\mathsf{P}(\varphi)$, the satisfaction relation $t\models\varphi$ is defined inductively by:
@@
For any set $\Gamma\cup\{\varphi\}$ of formulas, we write $t\models\Gamma$ if $t\models\psi$ for all $\psi\in\Gamma$, and we write $\Gamma\models\varphi$ if $t\models\Gamma$ implies $t\models\varphi$. We write $\varphi\models\psi$ for $\{\varphi\}\models\psi$. We write $\models\varphi$ if $t\models\varphi$ for all teams $t$ with domain containing $\mathsf{P}(\varphi)$.
$v_{pq}$$v_{p\overline{q}}$$v_{\overline{p}q}$$v_{\overline{pq}}$
(a) $t\models p\vee\lnot p$
$v_{pq}$$v_{p\overline{q}}$$v_{\overline{p}q}$$v_{\overline{pq}}$
(b) $s\models(p\land\textsc{ne})\vee(\lnot p\land\textsc{ne})$
$v_{pc}$$v_{p\overline{c}}$$v_{\overline{p}c}$$v_{\overline{pc}}$
(c) $u\models((p\land\textsc{ne})\vee(c\land\textsc{ne}))\land\textsc{ne}$$u\models\Diamond p\land\Diamond c$
Figure 1: Examples of the Semantics
Figure 1 displays some examples of the semantics. Each node represents a valuation, and is labelled according to what is evaluated as true ($1$) by that valuation, e.g., $v_{p\overline{q}}(p)=1$ and $v_{p\overline{q}}(q)=0$. The encircled area is a team we wish to highlight. The team $t$ in figure 1(a) satisfies the formula $p\vee\lnot p$ because it can be split into two subteams, one ($\{v_{p\overline{q}}\}$) satisfying the disjunct $p$, the other (the empty team $\emptyset$) satisfying the other disjunct $\lnot p$ (the connective $\vee$ is commonly known as the split disjunction). With the nonemptiness atom ne, we can enforce that teams be split into nonempty subteams: the team $s$ in Figure 1(b) satisfies the formula $(p\land\textsc{ne})\vee(\lnot p\land\textsc{ne})$ because it can be split into a nonempty team ($\{v_{pq}\}$) satisfying $p$, and a nonempty team ($\{v_{\overline{p}q}\}$) satisfying $\lnot p$.
Figure 1(c) relates to the applications of the atom ne in formal semantics and philosophical logic. On one possible interpretation of team semantics, teams represent information states: if the true information that one possesses is represented by a team $t$, one knows that the valuation that represents the way that things actually are (the actual world) is among the valuations in $t$, and hence whatever holds in all valuations in $t$ must also be the case in the actual world.
Consider the defined operator $\Diamond\varphi:=(\varphi\land\textsc{ne})\vee\top$—observe that the formula $\Diamond\varphi$ is true in a team $t$ just in case $t$ contains a nonempty subteam in which $\varphi$ is true. This operator is known as the might-operator: it, along with other similar connectives, has been used to model the meanings of epistemic possibility modalities such as the ‘might’ in “It might be raining” (see, e.g., [38, 40]). For if one’s information state contains a nonempty substate that supports the assertion that it is raining (if $t\models\Diamond r$), then the actual world might be one in which it is raining, and so, for all that one knows, it might be raining.
Now assume that Bob says, “Sue went to the park or to the cinema.” One natural inference from this might be that Bob deems it possible both that Sue went to the park and that she went to the cinema. The atom ne has been used to model inferences such as this one by Aloni and her collaborators (see, e.g., [4, 13]). On Aloni’s account, humans have a cognitive tendency to disregard structures that verify sentences by virtue of some empty configuration (the neglect-zero tendency). In the setting of team semantics, these empty configurations are represented by the empty team, and the neglect-zero tendency is modelled by appending $\land\textsc{ne}$’ to all subformulas of any formula that formalises a sentence that is being interpreted under this tendency. Then, for instance, the inference discussed above may be accounted for by pointing to the entailment $((p\land\textsc{ne})\vee(c\land\textsc{ne}))\land\textsc{ne}\models\Diamond p\land\Diamond c$—Bob’s utterance, when interpreted under the tendency (we append $\land\textsc{ne}$’ to all subformulas of the formula $p\vee c$, resulting in the formula $((p\land\textsc{ne})\vee(c\land\textsc{ne}))\land\textsc{ne}$) indicates that according to his information, Sue might have gone to the park and might have gone to the cinema ($\Diamond p\land\Diamond c$). See Figure 1(c). [^2] We have simplified and modified this account due to space constraints and to make it function in the propositional setting. For more on the actual account, see [4, 13].
@@
convexity property iff $[t\models\varphi,u\models\varphi,\text{ and }u\subseteq s\subseteq t]\implies s\models\varphi.$
It is easy to see that $\varphi$ is flat iff it has the empty team, downward closure, and union closure properties.
@@
The right-to-left direction follows by Proposition 1. For the left-to-right direction, an easy induction on the complexity of $\varphi$ shows that if ne occurs in $\varphi$, then $\varphi$ does not have the empty team property. We only give the disjunction case here. If $\varphi=\psi\vee\chi$ and ne occurs in $\varphi$, then it must occur in $\psi$ or $\chi$. Without loss of generality, assume that ne occurs in $\psi$. Then, by the induction hypothesis, $\psi$ does not have the empty team property, but then clearly neither does $\psi\vee\chi$.
@@ 3 Model-Checking Problem
Let $\mathrm{Maxsub}(t,\varphi)=\mathrm{maxsub}(t,\varphi)$ if $\mathrm{maxsub}(t,\varphi)$ exists, and $\mathrm{Maxsub}(t,\varphi)=\mathbf{0}$ otherwise, where $\mathbf{0}$ is a “null team” such that $\mathbf{0}\not\models\varphi$ for all $\varphi$. (Note that $\mathbf{0}$ is not an actual team, and in particular is not the empty team $\emptyset$—it is, rather, a notational shorthand we use to indicate that a formula $\varphi$ is not satisfiable in any subteam of a team $t$.) Let $|t|$ denote the size of $t$, and $|\varphi|$ the number of symbols in $\varphi$.
@@
First check whether $\varphi=\textsc{ne}$. If yes, then if $t=\emptyset$, return $\mathbf{0}$, and otherwise return $t$. Clearly this can be done in polynomial time with respect to $|t|+|\varphi|$. If $\varphi\neq\textsc{ne}$, check $\mathrm{maxsub}(t,\varphi)$; the result then follows by [22, Lemma 5.1].
@@
A function $f:\mathrm{subOcc}(\varphi)\to 2^{2^{\mathsf{P(\varphi)}}}\cup\{\mathbf{0}\}$ is called a labelling function. We describe an algorithm for computing a sequence $f_{0},f_{1},\ldots$ of such labelling functions. In what follows, we let $\mathbf{0}\cap s:=\mathbf{0}$ and $\mathbf{0}\cup s:=\mathbf{0}$ for all teams $s$.
@@ 5 Validity Problem
We make use of the following notion: the flattening $\varphi^{f}\in\mathrm{PL}$ of $\varphi\in\mathrm{PL}(\textsc{ne})$ is defined by $\varphi^{f}:=\varphi(\top/\textsc{ne})$ (see [6]; cf. [36]). It is easy to see that $\varphi\models\varphi^{f}$; observe also that by Proposition 2, $\varphi\in\mathrm{PL}(\textsc{ne})$ has the empty team property iff $\varphi=\varphi^{f}$.
@@ 6.0.1 Acknowledgements
The first author received support from the Academy of Finland, decision number 368671, and from the European Research Council (ERC), grant 101020762. The authors would like to thank Maria Aloni for helpful discussions related to the content of this article; as well as the anonymous reviewers, for their valuable comments and suggestions.